# The unified non-disciplinary model of the operation of neurons

### v3

@@ 3 Operation in a nutshell
The compression produces heat, as evidenced by a rise in temperature. The expansion consumes heat, as evidenced by a decrease in temperature. These are well-understood processes in thermodynamics. However, they are still unexplained in biophysics, although they were observed seven decades ago [22]. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). Correspondingly, one can observe, in addition to the electrical potential wave, mechanical, optical, and other features changing. In the discipline of electricity (when using the serial $RC$ model), it can be interpreted as a capacitive current, which physiology observes as hyperpolarization that inverts the polarity of the output voltage. (By mistake, physiology interprets the capacitive current that the outward current in the channel consists of $Na^{+}$ ions, and when the resultant potential changes its direction, the ion changes to $K^{+}$; a kind of alchemy.) In different disciplines, pressure and voltage represent the same phenomenon, using concepts from various disciplines. However, the speeds of those waves are strongly different: The AP as an electrical wave is a million times slower than the one described by the Maxwell laws, given that it is a side effect of the mechanical wave.
@@ 4 Membrane’s electricity
Biology represents a complex case where the phenomena cannot be reasonably reduced to a single discipline, as classic inanimate science does when deriving its laws for "material points", as A. Einstein coined, abstracted to either mass or charge, but not both. Moreover, the closed volume of the neurons drastically differs from the infinite space. The charges attract and repel each other, but since they experience the counterforce of the limiting surface (the membrane), a static mechanical pressure and a shock wave are also created, which naturally connect the electrical and mechanical effects, as well as the electrical, mechanical, and thermodynamic theories of operation. We must discuss them in a non-disciplinary way because the participating ions belong simultaneously to the disciplines of electricity and thermodynamics; furthermore, we must cross the borderline between particle-based and continuous points of view of nature [31]. Moreover, the features of living matter may change while being measured in physics laboratories as usual: "the construction is different from anything we have yet tested in the physical laboratory" [7].
@@ 4.1 Condenser
Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [32, 33], but does not explain that the negative and positive ends of lipids constituting the membrane, in an electrolytic solution, attract and absorb ions, forming a charged layer on the surfaces of the membrane. In this way, the membrane in electrolytes acts as a condenser. However, this condenser comprises two charged sheets on the membrane surfaces, with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (unlike conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolyte, so its thickness is far from the ideal zero thickness of classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electrical point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
@@ 4.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electrical forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electrical field (which can be of electrical, magnetic, mechanical, or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electrical field fluctuates, so their state is dynamic: molecules dissociate, and free ions recombine. Furthermore, the process also plays an important role in the energy supply of neurons: Adenozine Triphosphate (ATP) hydrolyzes molecules, and the electrical field near the membrane moves ions to the "condenser plates", thereby accumulating "potential energy". (The claim that the ions passing through ion channels do not use energy is wrong: they are accelerated by the condenser’s field and consume energy in that way.) Depending on their environment, substances can exist in base, ionized, and polarized states (and exhibit corresponding behaviors).
(For the sake of clarity, both mechanisms are known in physics. The term ’polarization’ is reserved for the redistribution of charges in multiply charged systems without the application of an external current. Changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even though, unfortunately, physiology refers to it as polarization. As discussed below, true polarization is also present, which confuses the description. External currents contribute to the changes of potential.)
@@ 4.3 Deriving the "thermodynamic electrical field"
known as Nernst equation. In other words, Eq.(1) and Eq.(2) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine quantities from the intracellular and extracellular sides additively, whether concentrations, mobilities, or Nernst voltages, as it happens in the Goldman-Hodgkin-Katz (GHK) equation (see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [31]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [6], the two differentials must change in the same direction, given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (2) results in a negative potential difference). (The derivation in ’ordinary physics’, such as  [34], Eq.(11.30) is wrong, at least because of two reasons: the same speed is assumed for mass and charge transport; furthermore, the used partial derivatives are not independent; the details are discussed throughout the paper.)
@@
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added a transport-related external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [36], with a lot of cited references.) This equation, expressing Newton’s first law, has no place for effects such as "Pumps that maintain ion gradients …transport ions against their electrical and chemical gradients[17], page 101. It implies some new force that moves ions, confronting Schrödinger’s science-based opinion: "And that not on the ground that there is any ‘new force’ or whatnot, directing the behaviour of the single atoms within a living organism"[7]. Biophysics still did not provide either the origin of the ’new force’ or the reason why Newton’s laws are not valid for ions in living matter, which undermines the credibility of those protein-based mechanisms.

### v4

@@ 3.2 Model operation
The pressure wave is well described in the thermodynamic view. The compression produces heat, as evidenced by a rise in temperature. The expansion consumes heat, as evidenced by a decrease in temperature. These are well-understood processes in thermodynamics. However, they are still unexplained in biophysics, although they were observed seven decades ago [26]. The elastic membrane functions as a damped oscillator, which, in its negative amplitude phase, reverses the direction of electrolyte flow (and, consequently, the direction of the current carried by the ions). Correspondingly, one can observe, in addition to the electrical potential wave, mechanical, optical, and other features changing. However, thermodynamics cannot account for any electrical change, given that electric charge is outside the region of its interest.
In the discipline of electricity, there is no direct way to interpret heat emission and absorption (see section 3.4.6), given that the heat is outside the notions used in electricity. However, considering the rush-in speed, one can give another reason and make a half-quantitative estimation of the amount of heat exchange.
3.3 Forces exerted on ions
Figure 1 shows the forces exerted on ions near the internal surface of the membrane at the exit of an ion channel in the moment next to that $Na^{+}$ ions rush into the internal segment of the neuron. The white (likely mainly $K^{+}$ with some $Na^{+}$) ions are sitting on the surface, and the charges on the opposite side of the membrane ($F_{Membrane}$) firmly fix them. They are not "free charge carriers": they provide the resting potential. They are at rest because of the counterforce exerted by the lipid layer.
At the beginning of the AP, $Na^{+}$ ions rush into the intracellular segment. The accelerating field across the membrane’s surfaces moves them at the corresponding Stokes-Einstein speed through the ion channel. When they reach the ion channel’s exit, in the absence of the enormous field, the ions slow down to the speed corresponding to the local field in the electrolyte; practically, in a short distance, they "stop". This way, a large amount (about $10^{7}$; about the amount of the free charge carriers in the segment) of ions appear suddenly in the layer near the surface; that is, the charge near the ion channel’s exit port suddenly increases. (In the previous moment, the ions present in the volume segment are free to move, but the resultant force of the thermodynamic and the electrical forces keeps them in place in a balanced state.) Since the total electric flux through any closed surface (Gaussian surface) is directly proportional to the enclosed electric charge, the local electric field suddenly increases compared to its environment. Hence, the ions attempt to move along the local gradient. There are constraints, however.
In the direction of the ion channel (where they arrived from). they face an electric field of opposite direction. In the resting state, most of the ions in that local volume are $K^{+}$ ions. The rushed-in ions suddenly increase the local $Na^{+}$ concentration that practically kicks out the $K^{+}$ ions, which help to block further ion input. Their appearance has two different effects in the directions perpendicular to and parallel with the membrane’s surface. In the perpendicular direction, relatively small $F_{Diff}$ forces move the ions toward the former ’resting state’ (the forces are composed of the local thermodynamic and electrical forces). Given that the fields are low, the speed of the in-place ions is low (much above the diffusion speed but much below the rush-in speed), a significant ion enrichment in the near-membrane region occurs. Notice that the ions move under the resultant force due to the thermodynamic and electrical gradients, so the $Na^{+}$ ions move toward the bulk, the $K^{+}$ ions toward the membrane (due to the different thermodynamic gradients). This movement, however, is slow compared to the speed of the gradient change of an AP process.
Figure 1: The forces action on ions at the moment when the rush-in finished.
In the horizontal direction, only the ions in the "dynamic layer" participate in the game. Three forces are exerted on the ions.
  • the increased electrical field pushes them toward the AIS
  • the increased local $Na^{+}$ concentration (thermodynamic plus electrical field) pushes the ions toward the AIS
  • the injected ions repel each other, so they arrive at the membrane’s wall, and they push the ions toward the AIS (a constraint force)
[…]
All this happens in the layer of thickness of 1 nm near the surface of the membrane, as observed by physiology. For the ions, the only way out is to leave the membrane through the AIS. The AIS, with its large number of ion channels, can be modeled as a resistor and generates a potential on it. As described, the potential changes its value and direction. Correspondingly, ions flow through the AIS initially initially outward, later inward, then again outward. There are no separated and cooperating $K^{+}$ and $Na^{+}$ channels with precisely timed activation. Instead, the mechanisms discussed above are valid. The AIS current consists mainly of $Na^{+}$ with a small amount of $K^{+}$. As discussed, new $Na^{+}$ ions from the outside space flow into the intracellular space. Although most of the $Na^{+}$ flows out through the AIS, the leftover diffuses into the bulk and must be pumped out. This fact is why Na/K pumps are needed.
At any given position, the Gauss flux of a volume continuously changes as the amount of charges in the given volume changes due to the vibration of the electrolyte, plus the mentioned forces. The electric field and the potential are proportional to the flux and so, to the charge, and so, to the current. The speed of the charge wave depends only on the medium (i.e., constant), and the amplitude is simply an exponential discharge. The observer can measure the effect.
Due to turbulence in the material flow, the "white" ions may also move toward the bulk for a short distance before reaching the dynamic layer. That is, all positive ions may move out through the AIS, without selectivity. As discussed in section 3.5, the hydrolyzis of Adenozine Triphosphate (ATP) produces "fresh" ions near the membrane and the local field moves them toward the membrane, so the neuron can restore its ’setpoint’ as described in section 5.2. That means the "downhill" movement of ions across the membrane through ion channels does not violate the conservation of energy. However, claims such as "Ion channels cannot be coupled to an energy source to perform active transport, so the transport that they mediate is always passive (’downhill’)" [27], are wrong. The surfaces of the membrane serve as a kind of ion- and energy buffer, and the continuous material transfer (the background logistics of life) performs refilling the buffer.
3.4 Connecting electricity and thermodynamics
The injected ions suddenly act on the elastic membrane as an "impulse force", so the membrane (and the slightly compressible electrolyte it contains) performs a damped oscillation. Given that the resultant of the above-mentioned electrical thermodynamic and constraint forces points toward the AIS, the ions will move with the corresponding Stokes-Einstein speed toward the AIS. However, $F_{Wall}$ is an oscillating force that changes its direction and may be larger than the other forces. The electrical and thermodynamic forces would essentially produce a simple discharge, with only one current direction. However, the damped oscillation has positive and negative phases, thus increasing and decreasing the electrical force. In a thermodynamic view, the membrane "sucks back" the electrolyte containing the ions. In the electrical view, the direction of the ion stream (aka current) changes to the opposite, causing the ’capacitive current’ to change the sign of the resultant current (the neuronal membrane can be modeled as a condenser). In a physiological view, polarization, depolarization, and hyperpolarization take place. As the model clearly suggests, all ions can be likely $K^{+}$ and $Na^{+}$ ions, the question is only their proportion.
$F_{Na^{+}}=10^{7}*1.60217663*10^{-19}[V][C]=1.6*10^{-12}\ [N]$ (1)
The pressure can be calculated using electrical parameters; i.e., in a cross-disciplinary way.
3.4.1 Mechanical consequences
[…]
Notice an important difference. The acceleration of an ion is unbelievably large. It is sufficiently large to keep the potential at the same value, provided that the ions must follow a small change, such as due to the "leaking" current through the AIS. Similarly, the ions can ’instantly’ follow quick changes such as a square wave gradient. However, if many similar ions are ahead, their repulsion decreases the acceleration, and the ion travels only at a few $m/s$ speed. The huge forces and accelerations mean that a potential change acts immediately. However, the force decays quickly. The ions start to move ’instantly’, but the charge carriers can move only with a limited speed, much below the interaction speed of electromagnetic (EM) interactions. The effect can propagate only with that lower speed. In the case of an axon, there exists a mechanical constraint that the ions cannot spread through the wall.
3.4.2 Speeds
A way to estimate the ions’ rush-in speed is that one assumes that the Stokes-Einstein speed describes the ions’ speed (Fig. 3 of [16], quantitatively underpins the hypothesis, see [34]), both in axons and ion channels. If so, the ratio of the potential gradients equals the ratio of the speeds of ions. One can estimate that the 100mV100~mV AP on the $50~\mu m$ AIS generates a $2*10^{4}~\frac{V}{m}$ electrical gradient and [16] measured $20~\frac{m}{s}$ for the speed of the AP. By assuming that the electrical field across the membrane is $10^{7}~\frac{V}{m}$, one can estimate the speed at the exit of the ion channels as $10^{4}~\frac{m}{s}$.
3.4.3 Energy relations
We can also estimate the work done when the rush-in charge extends the neuron, as the $P\times\Delta V$. We consider an average value $P$, and we calculate the change of the volume as neuron’s surface $10^{-8}\ m^{2}$ multiplied by the change of the radius $\Delta r=10^{-9}\ [m]$ multiplied by the pressure $P$, that gives $E_{mech}=2*10^{-14}\ [J]$. If one assumes that the pressure is of entirely mechanical origin [31], one must assume that the pressure is $10^{6}~\frac{N}{m^{2}}$, which is well above the measured value.
We can also estimate the electrical energy $E_{electr}=\frac{1}{2}C*(\Delta V)^{2}=1.4*10^{-12}\ [J]$, see also [35]. To estimate the time required to change the pressure, one can assume that the collision force of a $Na^{+}$ ion provides the pressure (the shock wave), and this can be calculated as $F=m\frac{\Delta v}{\Delta t}$. (Given that Eq. (1) provides the force, and one must assume that the ion loses its $10^{4}~to~10^{5}\frac{m}{s}$ speed, one concludes that it generates a shock wave in $\Delta t=10^{-8}~to~10^{-7}~seconds$.)
In line with the experimentally measured energy consumption [36] and our calculation in section 7.1, the AP cycle consumes energy in the range of $10^{-7}\ [J]$. The five orders of magnitude between the total and the electrical energy seems to underpin that in addition to the $10^{-3}$ concentration of ions, only a small fraction of the ions (the ones in the layers near the membrane) in the volume participate in the electric activity of issuing an AP, while the entire volume in the mechanical activity. The seven orders of magnitude between the total and the non-ionic thermodynamic energy suggest that classical thermodynamics is not valid for ions. That is, practically all the energy is invested as elastic energy, but the elasticity modulus of the membrane is exceptionally high. The primary reason for neuronal operation is the appearance of new ions in the vicinity of the membrane, but this creates a secondary, vast mechanical shock wave. By considering the electrical repulsion force between particles in a slow current, which is entirely omitted in the classic theory, we open the way for explaining the observed thermodynamic and mechanical changes. Instead of alternative disciplinary theories, cooperation between the classic disciplines is needed.
3.4.4 Electrical temperature
For thermodynamic distributions, one can interpret the "temperature of individual particles" with mean kinetic energy $E$ as $T=\frac{2*E}{3*k_{B}}$ (where $k_{B}$ is the Boltzmann constant and $T$ is the thermodynamic temperature of the bulk quantity of the substance); that is, the temperature is directly proportional to the average kinetic energy. The energy arrives at the neuron in the form of $10^{7}$ $Na^{+}$ ions, so the average energy a single ion conveys is $2.5*10^{-14}~[J]$. One can assume that most (99%) of that energy is consumed for moving the ion against the Stokes-Einstein force in the viscous medium, so after exiting the membrane, the speed of the ions is $\sqrt{\frac{2*2.5*10{-16}}{3.82*10^{-26}}}\approx 10^{5}~[m/s]$, which is well above the $\approx~500~m/s$ average speed.
The ions on the two sides of the membrane are in the state with temperature $T$ before and after generating the AP. When the $Na^{+}$ ions rush into the intracellular space, they gain energy through electrostatic acceleration by the membrane’s electrical potential, so the same ions appear on the intracellular side as having more energy, i.e., slightly higher temperature. The temperature is more than two orders of magnitude higher. However, the proportion of those "hot" ions is about six orders of magnitude lower, so one can expect a temperature change of up to a millikelvin, depending on the measuring conditions. In the second phase of the AP, those more energetic ions that provide excess local potential above the resting potential leave the membrane’s proximity through the AIS, so the temperature decreases in the second phase to its original value. Although it is hard to quantify the effect, it can contribute to the general heat production and adsorption (the thermodynamic interpretation of heat emission/absorption is described in section 3.4.6). In the original publication on solitons (essentially pressure waves) [17], a wide range of measured and theoretical data is discussed. They are in line with our conclusions, among others, $80\ \mu K$ measured temperature change. Furthermore, the electrical energy is about two orders of magnitude smaller than the elastic energy (it is called the soliton energy).
3.4.5 Resting vs. leakage current
Although through the thickness of the ion layer $\Delta z$, see section 4.4, the ion’s water-related behavior can influence the surface charge density, one can conclude that when only one type of ion is dissolved in the liquid, approximately the same concentration difference between the segments produces approximately the same "thermodynamic electric field" that can counterbalance the electric field of the biological condenser. One can notice that in the classic measurement by Hodgkin-Huxley, the sum of the concentrations of $K^{+}$ and $Na^{+}$ in the cytoplasm and extracellular fluid (in mM) are 380 and 390, respectively, while the sum concentrations of $Cl^{-}$ and $Ca^{2+}$ are 450 and 460; practically equal, as our analysis derived it must be. The deviations are well within the uncertainty of the measured values.
HH did a complete measurement, and derived precise measured data. However, their measurement was not accurate because their model was wrong. In a balanced state, no current flows, so no voltage is generated. For this reason, they assumed that a permanent leaking current $I_{L}$ was flowing on the membrane, and on its resistance, it generated a 42.5mV42.5\ mV voltage [37]. In HH’s picture, the question what ion constitutes the current remains open. It was claimed that the only permeability pathways open at rest are $g_{leak}$ and $g_{K}$. It follows that at the resting membrane potential, the leak current equals the potassium current. That is, at rest, two currents flow (without driving force in a balanced state), one consisting of $K^{+}$ ions and another of one $leak^{+/-}$ ions (??) (depending on the direction, it could be positive or negative), and neither of them changes the concentration neither on the departure nor on the arrival segment, neither the concentration of the other ions; defying also Nernst’s law.
The idea of leakage current was wrong: the voltage HH measured [37] is correct, but – as our model correctly explains – it was generated by charge separation instead of a current through a distributed resistor. It is permanently present while ionic segments are on its sides, even when no current flows. Measuring the energy consumption of neuronal operation [35] confirms that no resting current, in the sense as HH used the notion, exists.
3.4.6 Heat emission/absorption
There are alternative methods of producing a voltage difference across the membrane’s surface; with and without current. Hodgkin and Huxley made the wrong choice. The lack of resting current also solves the long-standing mystery of "reversible heat release during the action potential", which is inconsistent with the Hodgkin–Huxley model. The issue here is that the dissipation was attributed to the "leakage current." As our explanation shows, the voltage generated across the membrane results from charging rather than a permanently flowing current, so no dissipation occurs. Although the energy to build a potential on the membrane and to maintain it needs biological energy consumption, the process of generating an action potential is, in principle, mostly reversible: the energy is stored as elastic energy.
In the framework of the classic theory, Ohmic currents flow through resistors that dissipate heat due to friction, no matter in which direction ($W=I^{2}\times R$) the ion currents flow. The existence of "cooling", alone, undermines the credibility of the classic theory [38]. The issue here is that the dissipation was attributed to the "leakage current." As Hodgkin wrote: "Hill and his colleagues found [26] that an initial phase of heat liberation [of the action potential] was followed by one of heat absorption. […] a net cooling on open-circuit was totally unexpected and has so far received no satisfactory explanation." [39] "All authors came to similar conclusions: during the action potential, no significant net heat is produced. Transient heat releases are mostly reabsorbed in the second phase of the action potential. …The finding of a reversible heat release during the action potential of nerves is a striking and very fundamental fact. It is inconsistent with the HH model. The physics underlying the nervous impulse must rather be based on reversible processes". [40] For the correct thermodynamic explanation, see section 7.
The lack of leakage current (see section 3.4.5) not only validates our model, but also solves the long-standing mystery of reversible heat release during the AP. Pressure changes accompany the electrical process and naturally account for the observed temperature changes. Our discussion on the energy consumption of the nervous impulse underpins that the complex physical process comprises mostly reversible disciplinary processes: most of the energy of AP is stored as reversible elastic potential energy. Hence, its dissipation is also negligible: "no significant net heat is produced" [40]. (Alternatively, it may be hypothesized that the positive $K^{+}$ ions, while traveling ahead of the positive $Na^{+}$ ions under the effect of a potential-free space in the second phase of AP, not only turn with their positive charge the offset voltage to negative, but on the way the $K^{+}$ ions collect the heat dissipated by the $Na^{+}$ ions. "Laws of science cannot describe life".)
3.5 Background logistics
The neurons use a quasi-hidden background energy production and a state-restoration process that operates in parallel with neuronal function at a quasi-permanent speed. The ATP is produced and delivered to the spot by independent subsystems. The ATP’s delivery, similarly to that of the ions, obeys its component gradient. When the neuron’s membrane (through the AIS, resting ion channels, and pumps) ’loses’ ions, their local gradients in the vicinity of the membrane decrease. The interesting processes (ionization, ion absorption, and so on) occur near the membrane. The processes enable producing ’fresh’ ions from the available neutral molecules by hydrolysis to restore their concentration; then new ATP diffuses in the place of the ’used’ ones. Those ions build up on the membrane, increasing its electrical potential. That energy is a kind of potential energy and enables forwarding ions to the other segment (their presence slightly increases the potential there), while part of the energy is dissipated as the potential moves ions in the viscous fluid.
As discussed in section 4.5, outside the lipid bilayer, electric and concentration gradients exist. They enable the ions to move, but since their speeds are proportional to the resultant gradients, at different locations, the speeds of the macroscopic movement of ions, as well as the individual speeds of ions along their paths, can change by several orders of magnitude, from the drift speed to the potential-accelerated speed, while they travel a nanometer-long path under nanosecond periods. For example, a $Na^{+}$ ion is in rest (has the drift speed) on the high-concentration side when the ion channel opens, accelerates to a potential-accelerated speed, than slows down to the drift speed again on the low-concentration side. Even its velocity components in directions parallel and perpendicular to the membrane’s surface can differ by orders of magnitude when fields are exerted on them in the mentioned directions. The ion-gradient-related movement needs special care, since the driving force is different for the different chemical elements, and the movement of ions in close vicinity to each other (such as in ion channels) strongly affect each other.
@@ 4.1 Condenser
Cell biochemistry sufficiently well describes the lipid bilayers that constitute the membrane [49, 50]. The negative and positive ends of the lipids that constitute the membrane, in an electrolyte solution, attract and bind ions, forming a charged layer on the membrane surfaces. In this way, the membrane in electrolytes acts as a condenser, comprising two charged sheets on its surfaces, with ill-defined parameters. Their boundaries, thicknesses, and compositions depend on the operating state (unlike conducting plates with well-defined sizes and boundaries). This condenser is not ideal: the charged sheet comprises ions with attached aqueous constituents from the electrolyte, so its thickness is far from the ideal zero thickness of classical electricity. Furthermore, in its wall, it inseparably comprises ion channels (from an electrical point of view: resistors), and the condenser plates are in polarizable electrolyte solutions.
@@ 4.2 Electrolytes
Specific chemical substances can naturally hold either a positive or negative electrical charge and react to their micro-, macro-, and electrical/mechanical/thermodynamic environments. A molecule has internal electrical forces that keep its ions in place, so it has two charge centers (dipoles). When another dipole or a macroscopic external electrical field (which can be of electrical, magnetic, mechanical, or chemical origin) appears near the molecule, its perturbing effect can affect the relation of the ions to each other. The two charge centers initially increase their distance (the molecule polarizes). When that disturbance is strong enough, the ions can entirely separate (the molecule ionizes). The local electrical field fluctuates, so their state is dynamic: molecules dissociate, and free ions recombine. Furthermore, the process also plays an important role in the energy supply of neurons: ATP hydrolyzes molecules, and the electrical field near the membrane moves ions to the "condenser plates", thereby accumulating "potential energy". (The claim that the ions pass through ion channels by using the "downhill method" does not use energy, is wrong: they are accelerated by the condenser’s field and consume energy in that way.) Depending on their environment, substances can exist in base, ionized, and polarized states (and exhibit corresponding behaviors).
(In physics, the term ’polarization’ is reserved for the redistribution of charges in multiply charged systems without generating an external current. Changes in the membrane’s potential difference are a consequence of complete charge separation (and the movement of slow charges), even though, unfortunately, physiology refers to it as polarization. As discussed below, true polarization is also present, which confuses the description. External currents contribute to the changes of potential.)
@@ 4.3 Deriving the "thermodynamic electrical field"
known as Nernst equation. In other words, Eq.(3) and Eq.(4) are the differential and integral formulations of the same knowledge. The limits of the integration are chosen arbitrarily (by choosing the reference concentration $C_{k}^{int}$). Hence, the derived potential inherently comprises an additive term (a potential difference), so they must not be compared directly if they use a different reference $C_{k}^{int}$ (furthermore, it is nonsense to combine quantities from the intracellular and extracellular sides additively, whether concentrations, mobilities, or Nernst voltages, as it happens in the Goldman-Hodgkin-Katz (GHK) equation; see section 5.3). The derivatives are spatial derivatives; the temporal derivatives (needed for describing the time course) of the concentration(s) and voltage are derived in [48]. In ’ordinary’ physics, where the charge and mass are independent, changing one side changes the other in the opposite direction. In ’non-ordinary’ physics [8], the two differentials must change in the same direction, given that ions’ charge and mass cannot be separated (BTW: this is why Eq. (4) results in a negative potential difference). (The derivation in ’ordinary physics’, such as  [51], Eq.(11.30) is wrong, at least because of two reasons: the same speed is assumed for mass and charge transport; furthermore, the used partial derivatives are not independent; the details are discussed throughout the paper.)
@@
We multiplied the usual two terms by the elementary charge, so its terms are expressed as forces, plus we added a transport-related external force (its role is discussed below). Furthermore, changing one force triggers a corresponding counterforce and/or causes the ion to move in a viscous fluid. (An excellent summary of the possible biology-related forces is given in [53], with a lot of cited references.) This equation, expressing Newton’s first law, has no place for effects such as "Pumps that maintain ion gradients …transport ions against their electrical and chemical gradients[19], page 101. See also section 4.10.5. Biophysics still did not provide either the origin of the ’new force’ or the reason why Newton’s laws are not valid for ions in living matter, which undermines the credibility of those protein-based mechanisms.
@@
We call attention to the statistical nature of the derived thermodynamic force. It is valid as long as the conditions of the Nernst-Planck equation are valid; furthermore, it can be used with the same limitations. That is, one cannot use it for describing the movement of individual ions, but it can describe the charge/mass transfer represented by a statistical population of ions.
An interesting option is when the counterforce combines two disciplines, see section 3.4. When $Na^{+}$ ions rush-in at the beginning into the membrane, they produce a huge electrical and thermodynamic gradient simultaneously. The repulsion among the ions creates a huge mechanical pressure that presses the elastic membrane. The counterforce starts a mechanical shock wave (soliton) [17], that is simultaneously an electrical potential wave [18] that is measured as AP. Only cross-disciplinary discussion (non-ordinary laws) enables understanding of neuronal operation. That means when describing an ionic transfer process, we must not separate the electrical current from the mass transfer: they happen simultaneously and mutually trigger each other. Notice that the thermodynamic term is ion-specific while the electrical term is not. To be entirely balanced, the system must be balanced to all elements. In this way, changing one concentration implicitly changes all other concentrations and the electrical field.
It was a colossal mistake to introduce equivalent circuits with their fixed-voltage generators. It forces one to assume that the conductances of the players (membrane, synapses, AIS) change without any reason, and prevents understanding how the competition between thermodynamic and electrical processes governs neuronal operation. It leads, among others, to attributing conductance change to membranes, which are simple isolators with no charge carriers that can implement charge transfer, see section 4.8; this way attributing the change of an electrical entity to the biological material. This assumption neglects the driving force that the mentioned forces provide. Instead, it attributes the magic ability to the ion channels that they can change their transmission ability as the actual situation requires.
In balanced states, no transport occurs, so the transport force cancels, and the mobility has no role. The counterforce adapts to the situation. That force may be a mechanical one: the ions sitting on the surface of the membrane press the surface due to the attractive force of ions, and the membrane mechanically provides the needed counterforce. If the boundary of the segments is not freely penetrable, the counterforce equals the difference between those two forces. In this way, no force acts on the ions; the two gradients persist. When ions can move freely between segments, they will continue to move until they establish a concentration gradient (a thermodynamic force) for that ion.